The following several questions will be based on this example. In a small town there are...

  1. The following several questions will be based on this example.

    In a small town there are two doctor's offices (A and B). Each is considering acquiring a new piece of medical equipment that costs $50,000. Total demand in the small town and surrounding area for this new equipment amounts to $75,000. Each firm has to decide whether to get the new technology, and the new patients/business that it will bring in. If both get the new equipment, then they're each paying the full cost for the equipment, and each will get half of the new business. If only one gets the equipment, they still pay the full cost but get all the new business; the other practice would lose $10,000 in business because patients will go to the practice with new equipment. If neither buys, patients do not get access to the new equipment, no costs are incurred, and no new business is generated.

    Assume practices make a decision without knowing the decision of the other practice. What is the rational choice for Practice A, and why?

    Get equipment; maximizes income regardless what B does

    Don't get equipment; it won't help anyone

    Get equipment; profit is guaranteed

    No way to know

What about Practice B, assuming they have to decide without knowing A's choice?

Get equipment; maximizes income regardless what A does

No way to know

Get equipment; profit is guaranteed

Don't get equipment; it won't help anyone
Answer #1

Assume all the four cases payoff for the situation A buy B buy each will face loss of 75000/2-50000=-12500 each,A buy B not buy profit for A=75000-50000=25000 loss for B =-10000, B buy A not buy, profit for B =25000, loss for A=-10000, both do not buy 0 profit for each, Since there is no dominant strategy for any of the person as buying may give you profit of 25000 or loss of -12500 same way not buying may give you 0 profit or loss of -10000 there is no dominant strategy hence the correct answer would be No way to know for either of them